# TiRex: Zero-Shot Forecasting Across Long and Short Horizons with Enhanced In-Context Learning

### v1

@@ 1 Introduction
Figure 1: Two exemplary time series from the GiftEval benchmark. For both examples, we show one plot with the full context and TiRex’s prediction, as well as zoomed-in forecasts of the best-performing zero-shot models. Each plot shows the signal in blue, the model’s (median) prediction in orange, and the uncertainty bounds in gray. (a) A time series that exhibits strong peaks. Only TiRex is capable of predicting the periodic short spikes. (b) A time series with strong but noisy periodical behavior. TiRex predicts a meaningful uncertainty estimate (quantile range) over the long forecast horizon, while TimesFM and Chronos Bolt struggle because of collapsing quantiles.
@@ 1.2 Related Work
Currently, pre-trained time series models, with the capability of zero-shot generalization across datasets, predominantly adopt different transformer architectures. For instance, Chronos (Ansari et al., 2024a) and Chronos-Bolt (Ansari et al., 2024b) use an encoder-decoder variant. Moirai (Woo et al., 2024) adopts an encoder-only design with a masked modeling objective, and TimesFM (Das et al., 2024) follows a decoder-only causal modeling strategy for autoregressive generation. TabPFN (Hollmann et al., 2025) and its adaptation to time series TabPFN-TS (Hollmann et al., 2025) use a modified transformer encoder and pre-train only on synthetic data. A notable exception is TTM (Ekambaram et al., 2024), since it builds on the MLP-based TSMixer architecture (Chen et al., 2023).
@@ 3 Data Augmentation
To facilitate more diverse time series patterns and enhance the model’s exposure to a wider range of potentially relevant dynamics, we propose three augmentations for pre-training. This is inspired by the successful application of augmentation techniques in pre-training of other modalities, e.g., vision (Tian et al., 2020). The employed augmentations, illustrated in Figure 3, are: (1) Amplitude Modulation, which introduces trends and change points in the scale of the time series. Formally, a time series $\mathbf{y}$ is transformed by $y^{\prime}_{t}=y_{t}\cdot a_{t}$, where $a_{t}$ follows a linear trend (potentially with change points). (2) Censor Augmentation, which censors values within the time series at a random threshold. The augmented respective series is computed by $y^{\prime}_{t}=\text{max/min}(y_{t},c)$, The censor threshold $c$ is sampled by uniformly drawing a quantile from the empirical distribution of the signal. (3) Spike Injection, which adds short, periodic spike signals to the time series. The augmented time series is computed by $y^{\prime}_{t}=y_{t}+s_{t}$, where $s_{t}$ represents the added spike signals. The shape of each spike is defined by a kernel, which can be a tophat, a radial basis function (RBF), or a linear kernel. The periodicity and the specific parameters of the kernel (e.g., width, height for tophat; center, variance for RBF) are sampled from predefined distributions for each sample. The augmentations are applied with a probability of $0.5$ for amplitude modulation and censor augmentation and $0.05$ for spike augmentation. Appendix B provides a more detailed description of the augmentation procedures.
@@ 4 Experiments
The evaluation follows the respective benchmark protocols using mean absolute scaled error (MASE) for point forecast performance and the continuous ranked probability score (CRPS) for probabilistic forecast performance. Practically, the CRPS is approximated by the mean weighted quantile loss (WQL) over nine quantiles: 0.1 to 0.9 in increments of 0.1333In the Chronos-ZS benchmark, the mean weighted quantile loss (WQL) is used directly, hence, essentially both benchmarks evaluate on the same metric. Aggregated performance is computed by normalizing each evaluation setting’s score by that of a seasonal naive baseline, followed by the geometric mean across evaluation settings. Additionally, the average rank across evaluation settings is reported to ensure robustness against outlier performance.
@@ 4.1 Zero-Shot Forecasting
In this benchmark TiRex consistently outperforms all competing methods across both short- and long-term forecasting tasks (Figure 4). In terms of CRPS, TiRex achieves a score of $0.411$ (with standard deviation over training with 6 seeds of ±$\pm 0.002$), notably surpassing the next best zero-shot models ($0.459$, $0.463$, $0.481$). This performance gap is also reflected in the average rank, where TiRex shows a substantial lead over the second-best model, while the three models that following in the ranking (TimesFM-2.0, TabPFN, and Chronos-Bolt-Base) achieve very similar scores among themselves. Importantly, while other models tend to peak either in short- (e.g., Chronos-Bolt) or long-term (e.g., TabPFN-TS) forecasting, TiRex is the only model to excel simultaneously at both. Moreover, TiRex attains these results with significantly fewer parameters (35M) compared to Chronos-Bolt-Base (200M) and TimesFM-2.0 (500M). The advantage is most pronounced in long-term forecasting, where TiRex becomes the first zero-shot model to surpass the performance of PatchTST and TFT.
@@
Apart from the forecasting performance, we analyze the GPU memory consumption and inference runtime across all models. Specifically, we evaluate samples with a context length of $2048$ and a prediction length of $32$ over multiple batch sizes. As expected, given TiRex substantially smaller size compared to the next best models (TimesFM-2.0 and Chronos-Bolt-Base), TiRex requires significantly less GPU memory and achieves faster inference speeds. Specifically, TiRex is over $\times1$ faster than TimesFM-2.0, over 4×4$\times$4 × faster than Chronos-Bolt Base, and over 2176×2176$\times$2176 × faster than TabPFN-TS. Furthermore, TiRex even outperforms Chronos-Bolt Small, a similarly sized transformer-based architecture for larger batch sizes. The differences in maximum GPU memory consumption follow a similar order.
@@ 4.2 Ablations
We analyze the effectiveness of the proposed Contiguous Patch Masking (CPM) procedure by comparing three configurations: (1) standard decoder-only training with autoregressive inference — this setting is similar e.g., to TimesFM; (2) training with a naive multiple-patch procedure where we place the “rollout” always at the end of the sequence; and (3) training with CPM. Configurations (2) and (3) use the same inference procedure with masked tokens for multi-step forecasting, whereas (1) relies on patch-wise autoregressive decoding. As shown in Table 4.2 only CPM enables strong long-term performance without degrading short-term performance. In contrast, naïve multi-patch training diminishes the short-term forecasting performance, and standard next token training combined with autoregressive inference harms the long-term forecasting performance. These findings suggest that CPM is essential for training and inference behaviors under multi-patch prediction.
@@
To assess the impact of our augmentations, we trained our model excluding each augmentation and with no augmentations at all. The results, detailed in Table 4.2, indicate that including the augmentations is beneficial, i.e., improves the performance of the model. Performance consistently decreased in at least one benchmark metric when any single augmentation was removed. The most substantial decline occurs when no augmentations were applied. This underscores the effectiveness of each augmentation and their combined positive impact on the model’s generalization capabilities. Appendix D.3 provides a sensitivity analysis in terms of application probability.
@@
To assess the architectural choice of xLSTM with sLSTM modules, we replace it with mLSTM (Beck et al., 2024) and transformer blocks (Touvron et al., 2023) while keeping the patching and training procedure unchanged. For the transformer variant, rotary positional embeddings (Su et al., 2024) are added to mitigate the absence of inherent positional information, due to their permutation-invariance property. We also analyze mLSTM and sLSTM mix architectures as proposed in the original xLSTM paper. We denote xLSTM[i:j] for an architecture where $i$ sLSTM blocks are combined with $j$ mLSTM blocks. Additionally, we ablate our overall architecture by comparing it to a Chronos-Bolt architecture (Base and Small) that we train with our training procedure, using the same datasets and augmentations as for TiRex. Table 4.2 summarizes the results. TiRex with only sLSTM blocks yields the best performance, especially on long-term forecasts. We hypothesize that this is because its explicit state-tracking capabilities (Beck et al., 2024), which might facilitate uncertainty propagation and enable accurate modeling of periodic temporal structures over extended horizons. Using only mLSTM performs worst. However, switching just one of these blocks back to a sLSTM improves the results close to the sLSTM-only architecture of TiRex. The comparison to a Chronos Bolt architecture, which is consistently outperformed by TiRex, highlights that our overall architecture is critical to achieve good performance on both long and short-term forecasting. A more detailed comparison to the Chronos Bolt architecture is presented in Appendix D.3.
Table 1: Ablation study of individual components. The top two rows report the mean and standard deviation of TiRex over six runs with different random seeds. For the ablation variants, results that degrade performance by more than 3$\times$ the standard deviation relative to TiRex are underlined. Columns correspond to evaluation settings: GiftEval-ZS benchmark (overall, short-term, and long-term) and Chronos-ZS benchmark. Lower values indicate better performance.
@@
Augment w/o any $\underline{0.430}$ $\underline{0.682}$ $\underline{0.339}$ $\underline{0.478}$ $\underline{0.473}$ $\underline{0.722}$ $\underline{0.623}$ $\underline{0.800}$
@@
Backbone Transformer $\underline{0.422}$ $\underline{0.662}$ $\underline{0.342}$ $\underline{0.472}$ $\underline{0.461}$ $0.702$ $0.597$ $0.768$
@@
[…]
@@ 5 Conclusion
Like most pre-trained forecasting models, TiRex focuses on univariate time series. Although modeling multivariate series as independent univariate signals often performs well — as reflected in the GiftEval results and, for example, shown in Nie et al. (2022) — future work could incorporate multivariate data, for example in the form of extended contexts or modified input layers. Due to computational constraints, we did not extensively tune hyperparameters and only conducted a sensitivity analysis on key parameters. Future work should explore more comprehensive tuning for additional performance gains.
@@
We thank Maximilian Beck and Korbinian Pöppel for the discussions and advice in regard to xLSTM. The ELLIS Unit Linz, the LIT AI Lab, and the Institute for Machine Learning are supported by the Federal State Upper Austria.
@@
xLSTM introduces two cell types: the matrix LSTM (mLSTM) and the scalar LSTM (sLSTM). The mLSTM is designed to increase memory capacity through a matrix-based memory representation, and enables efficient parallel computation. In contrast, the sLSTM preserves a true recurrent pathway as in the original LSTM, enabling strong state-tracking capabilities (Beck et al., 2024). The recurrent pathway makes LSTM more expressive than State Space Models (SSMs), parallelizable RNNs like RWKV, and transformers (Merrill & Sabharwal, 2023; Merrill et al., 2024; Delétang et al., 2023). Figure 7 illustrates the respective expressivity hierarchies. Delétang et al. (2023) demonstrate on synthetic language tasks that this state-tracking capability of LSTMs yields empirical advantages. Beck et al. (2024) show that sLSTM retains these advantages. sLSTM employs a multi-head strategy along the recurrent pathway to improve efficiency.
@@
$\bm{c}_{t}$ $=\ \mathbf{f}_{t}\odot\bm{c}_{t-1}\ +\ \mathbf{i}_{t}\odot\bm{z}_{t}\ ,$ cell state (3)
$\bm{n}_{t}$ $=\ \mathbf{f}_{t}\odot\bm{n}_{t-1}\ +\ \mathbf{i}_{t}\ ,$ normalizer state (4)
$\bm{h}_{t}$ $=\ \mathbf{o}_{t}\odot\tilde{\bm{h}}_{t}\ ,$ $\tilde{\bm{h}}_{t}$ $=\ \bm{c}_{t}\odot\bm{n}_{t}^{-1}$ hidden state (5)
$\bm{z}_{t}$ $=\ \varphi\left(\tilde{\bm{z}}_{t}\right)\ ,$ $\tilde{\bm{z}}_{t}$ $=\ \bm{W}_{\bm{z}}\ \bm{x}_{t}\ +\ \bm{R}_{\bm{z}}\ \bm{h}_{t-1}\ +\ \bm{b}_{\bm{z}}$ cell input (6)
$\mathbf{i}_{t}$ $=\ \exp\left(\tilde{\mathbf{i}}_{t}\right)\ ,$ $\tilde{\mathbf{i}}_{t}$ $=\ \bm{W}_{\mathbf{i}}\ \bm{x}_{t}\ +\ \bm{R}_{\mathbf{i}}\ \bm{h}_{t-1}\ +\ \bm{b}_{\mathbf{i}}$ input gate (7)
$\mathbf{f}_{t}$ $=\ \exp\left(\tilde{\mathbf{f}}_{t}\right)\ ,$ $\tilde{\mathbf{f}}_{t}$ $=\ \bm{W}_{\mathbf{f}}\ \bm{x}_{t}\ +\ \bm{R}_{\mathbf{f}}\ \bm{h}_{t-1}\ +\ \bm{b}_{\mathbf{f}}$ forget gate (8)
[…]
@@
Context-free (CF) Push-down automaton (PDA) $+$ infinite stack
Context-sensitive (CS) Linear bounded automaton (LBA) $+$ bounded tape
Recursively enumerable (RE) Turing machine (TM) $+$ infinite tape

### v2

@@ 1 Introduction
Figure 1: Two exemplary time series from the GiftEval benchmark. For both examples, we show one plot with the full context and TiRex’s prediction, as well as zoomed-in forecasts of the best-performing zero-shot models. Each plot shows the ground truth signal in blue, the model’s (median) prediction in orange, and the uncertainty bounds in gray. (a) A time series that exhibits strong peaks. Only TiRex is capable of predicting the periodic short spikes. (b) A time series with strong but noisy periodical behavior. TiRex predicts a meaningful uncertainty estimate (quantile range) over the long forecast horizon, while TimesFM and Chronos Bolt struggle because of collapsing quantiles.
@@ 1.2 Related Work
Currently, pre-trained time series models, with the capability of zero-shot generalization across datasets, predominantly adopt different transformer architectures. For instance, Chronos (Ansari et al., 2024a), Chronos-Bolt (Ansari et al., 2024b), and COSMIC (Auer et al., 2025b) use an encoder-decoder variant. Moirai (Woo et al., 2024) adopts an encoder-only design with a masked modeling objective, and TimesFM (Das et al., 2024) follows a decoder-only causal modeling strategy for autoregressive generation. TabPFN (Hollmann et al., 2025) and its adaptation to time series TabPFN-TS (Hoo et al., 2025) use a modified transformer encoder and pre-train only on synthetic data. A notable exception is TTM (Ekambaram et al., 2024), since it builds on the MLP-based TSMixer architecture (Chen et al., 2023).
@@ 3 Data Augmentation
To facilitate more diverse time series patterns and enhance the model’s exposure to a wider range of potentially relevant dynamics, we propose three augmentations for pre-training. This is inspired by the successful application of augmentation techniques in pre-training of other modalities, e.g., vision (Tian et al., 2020). The employed augmentations, illustrated in Figure 3, are: (1) Amplitude Modulation, which introduces trends and change points in the scale of the time series. Formally, a time series $\mathbf{y}$ is transformed by $y^{\prime}_{t}=y_{t}\cdot a_{t}$, where $a_{t}$ follows a linear trend (potentially with change points). (2) Censor Augmentation, which censors values within the time series at a random threshold. The augmented respective series is computed by $y^{\prime}_{t}=\text{max/min}(y_{t},c)$, The censor threshold $c$ is sampled by uniformly drawing a quantile from the empirical distribution of the signal. (3) Spike Injection, which adds short, periodic spike signals to the time series. The augmented time series is computed by $y^{\prime}_{t}=y_{t}+s_{t}$, where $s_{t}$ represents the added spike signals. The shape of each spike is defined by a kernel, which can be a tophat, a radial basis function (RBF), or a linear kernel. The periodicity and the specific parameters of the kernel (e.g., width, height for tophat; center, variance for RBF) are sampled from predefined distributions for each sample. This randomization is designed to encourage the model to learn a general concept of transient events, rather than memorizing specific periodic patterns. The augmentations are applied with a probability of $0.5$ for amplitude modulation and censor augmentation and $0.05$ for spike augmentation. Appendix B provides a more detailed description of the augmentation procedures.
@@ 4 Experiments
The evaluation follows the respective benchmark protocols using mean absolute scaled error (MASE) for point forecast performance and the continuous ranked probability score (CRPS) for probabilistic forecast performance. Practically, the CRPS is approximated by the mean weighted quantile loss (WQL) over nine quantiles: 0.1 to 0.9 in increments of 0.1333In the Chronos-ZS benchmark, the mean weighted quantile loss (WQL) is used directly, hence, essentially both benchmarks evaluate on the same metric. Aggregated performance is computed by normalizing each evaluation setting’s score by that of a seasonal naive baseline, followed by the geometric mean across evaluation settings [^4] GiftEval benchmark scores are reported based on the leaderboard computation at the time of our submission. A subsequent update to the seasonal naive baseline affects the absolute aggregated scores but does not change any discussed model ranking, result, or conclusion.. Additionally, the average rank across evaluation settings is reported to ensure robustness against outlier performance.
@@ 4.1 Zero-Shot Forecasting
In this benchmark TiRex consistently outperforms all competing methods across both short- and long-term forecasting tasks (Figure 4). In terms of CRPS, TiRex achieves a score of $0.411$ (with standard deviation over training with 6 seeds of $\pm 0.002$), notably surpassing the next best zero-shot models ($0.459$, $0.463$, $0.481$). This performance gap is also reflected in the average rank, where TiRex shows a substantial lead over the second-best model, while the three models that following in the ranking (TimesFM-2.0, TabPFN, and Chronos-Bolt-Base) achieve very similar scores among themselves. Importantly, while other models tend to peak either in short- (e.g., Chronos-Bolt) or long-term (e.g., TabPFN-TS) forecasting, TiRex is the only model to excel simultaneously at both. Moreover, TiRex attains these results with significantly fewer parameters (35M) compared to Chronos-Bolt-Base (200M) and TimesFM-2.0 (500M). The advantage is most pronounced in long-term forecasting, where TiRex becomes the first zero-shot model to surpass the performance of PatchTST and TFT.
@@
Apart from the forecasting performance, we analyze the GPU memory consumption and inference runtime across all models. Specifically, we evaluate samples with a context length of $2048$ and a prediction length of $32$ over multiple batch sizes. As expected, given TiRex substantially smaller size compared to the next best models (TimesFM-2.0 and Chronos-Bolt-Base), TiRex requires significantly less GPU memory and achieves faster inference speeds. Specifically, TiRex is over $11\times$ faster than TimesFM-2.0, over $4\times$ faster than Chronos-Bolt Base, and over $2176\times$ faster than TabPFN-TS. Furthermore, TiRex even outperforms Chronos-Bolt Small, a similarly sized transformer-based architecture for larger batch sizes. The differences in maximum GPU memory consumption follow a similar order.
@@ 4.2 Ablations
We analyze the effectiveness of the proposed Contiguous Patch Masking (CPM) procedure by comparing three configurations: (1) standard decoder-only training with autoregressive inference — this setting is similar e.g., to TimesFM; (2) training with a naive multiple-patch procedure where we place the “rollout” always at the end of the sequence; and (3) training with CPM. Configurations (2) and (3) use the same inference procedure with masked tokens for multi-step forecasting, whereas (1) relies on patch-wise autoregressive decoding. As shown in Table 1 only CPM enables strong long-term performance without degrading short-term performance. In contrast, naïve multi-patch training diminishes the short-term forecasting performance, and standard next token training combined with autoregressive inference harms the long-term forecasting performance. These findings suggest that CPM is essential for training and inference behaviors under multi-patch prediction.
@@
To assess the impact of our augmentations, we trained our model excluding each augmentation and with no augmentations at all. The results, detailed in Table 1, indicate that including the augmentations is beneficial, i.e., improves the performance of the model. Performance consistently decreased in at least one benchmark metric when any single augmentation was removed. The most substantial decline occurs when no augmentations were applied. This underscores the effectiveness of each augmentation and their combined positive impact on the model’s generalization capabilities. Appendix D.4 provides a sensitivity analysis in terms of application probability.
@@
To assess the architectural choice of xLSTM with sLSTM modules, we replace it with mLSTM (Beck et al., 2024) and transformer blocks (Touvron et al., 2023) while keeping the patching and training procedure unchanged. For the transformer variant, rotary positional embeddings (Su et al., 2024) are added to mitigate the absence of inherent positional information, due to their permutation-invariance property. We also analyze mLSTM and sLSTM mix architectures as proposed in the original xLSTM paper. We denote xLSTM[i:j] for an architecture where $i$ sLSTM blocks are combined with $j$ mLSTM blocks. Additionally, we ablate our overall architecture by comparing it to a Chronos-Bolt architecture (Base and Small) that we train with our training procedure, using the same datasets and augmentations as for TiRex. Table 1 summarizes the results. TiRex with only sLSTM blocks yields the best performance, especially on long-term forecasts. We hypothesize that this is because its explicit state-tracking capabilities (Beck et al., 2024), which might facilitate uncertainty propagation and enable accurate modeling of periodic temporal structures over extended horizons. Using only mLSTM performs worst. However, switching just one of these blocks back to a sLSTM improves the results close to the sLSTM-only architecture of TiRex. The comparison to a Chronos Bolt architecture, which is consistently outperformed by TiRex, highlights that our overall architecture is critical to achieve good performance on both long and short-term forecasting. A more detailed comparison to the Chronos Bolt architecture is presented in Appendix D.4.
[…]
@@
Augment
w/o any $\underline{0.430}$ $\underline{0.682}$ $\underline{0.339}$ $\underline{0.478}$ $\underline{0.473}$ $\underline{0.722}$ $\underline{0.623}$ $\underline{0.800}$
@@
Backbone
Transformer $\underline{0.422}$ $\underline{0.662}$ $\underline{0.342}$ $\underline{0.472}$ $\underline{0.461}$ $0.702$ $0.597$ $0.768$
@@
Chronos Bolt S $\underline{0.456}$ $\underline{0.676}$ $\underline{0.413}$ $\underline{0.498}$ $\underline{0.463}$ $0.705$ $0.609$ $\underline{0.791}$
@@ 5 Conclusion
Like most pre-trained forecasting models, TiRex focuses on univariate time series. Although modeling multivariate series as independent univariate signals often performs well — as reflected in the GiftEval results and, for example, shown in Nie et al. (2022) — future work could incorporate multivariate data, for example in the form of extended contexts or modified input layers. Due to computational constraints, we did not extensively tune hyperparameters and only conducted a sensitivity analysis on key parameters. Future work should explore more comprehensive tuning for additional performance gains and investigate leveraging the model’s learned representations for other downstream tasks, such as classification (Auer et al., 2025a) or anomaly detection.
@@
We thank Maximilian Beck and Korbinian Pöppel for the discussions and advice with regard to xLSTM. We thank Elias Bürger, Bernhard Voggenberger, Marco Obermeier, and Levente Zolyomi for their efforts in providing an updated TiRex 1.1. We thank Martin Loretz for making TiRex run efficiently on CPU. The ELLIS Unit Linz, the LIT AI Lab, and the Institute for Machine Learning are supported by the Federal State Upper Austria.
@@
xLSTM introduces two cell types: the matrix LSTM (mLSTM) and the scalar LSTM (sLSTM). The mLSTM is designed to increase memory capacity through a matrix-based memory representation, and enables efficient parallel computation. In contrast, the sLSTM preserves a true recurrent pathway as in the original LSTM, enabling strong state-tracking capabilities (Beck et al., 2024). The recurrent pathway makes LSTM more expressive than State Space Models (SSMs), parallelizable RNNs like RWKV, and transformers (Merrill & Sabharwal, 2023; Merrill et al., 2024; Delétang et al., 2023). Figure 7 illustrates the respective expressivity hierarchies. We hypothesize that this advantage in expressivity allows TiRex to better model complex temporal dynamics, leading to improved forecasting performance, especially over long horizons. Delétang et al. (2023) demonstrate on synthetic language tasks that this state-tracking capability of LSTMs yields empirical advantages. Beck et al. (2024) show that sLSTM retains these advantages. sLSTM employs a multi-head strategy along the recurrent pathway to improve efficiency.
@@
$\bm{c}_{t}\$$=\ \mathbf{f}_{t}\odot\bm{c}_{t-1}\ +\ \mathbf{i}_{t}\odot\bm{z}_{t}\ ,$ cell state (3)
$\bm{n}_{t}\$$=\ \mathbf{f}_{t}\odot\bm{n}_{t-1}\ +\ \mathbf{i}_{t}\ ,$ normalizer state (4)
$\bm{h}_{t}\$$=\ \mathbf{o}_{t}\odot\tilde{\bm{h}}_{t}\ ,$ $\tilde{\bm{h}}_{t}\$$=\ \bm{c}_{t}\odot\bm{n}_{t}^{-1}$ hidden state (5)
$\bm{z}_{t}\$$=\ \varphi\left(\tilde{\bm{z}}_{t}\right)\ ,$ $\tilde{\bm{z}}_{t}\$$=\ \bm{W}_{\bm{z}}\ \bm{x}_{t}\ +\ \bm{R}_{\bm{z}}\ \bm{h}_{t-1}\ +\ \bm{b}_{\bm{z}}\ \$cell input (6)
$\mathbf{i}_{t}\$$=\ \exp\left(\tilde{\mathbf{i}}_{t}\right)\ ,$ $\tilde{\mathbf{i}}_{t}\$$=\ \bm{W}_{\mathbf{i}}\ \bm{x}_{t}\ +\ \bm{R}_{\mathbf{i}}\ \bm{h}_{t-1}\ +\ \bm{b}_{\mathbf{i}}\ \$input gate (7)
$\mathbf{f}_{t}\$$=\ \exp\left(\tilde{\mathbf{f}}_{t}\right)\ ,$ $\tilde{\mathbf{f}}_{t}\$$=\ \bm{W}_{\mathbf{f}}\ \bm{x}_{t}\ +\ \bm{R}_{\mathbf{f}}\ \bm{h}_{t-1}\ +\ \bm{b}_{\mathbf{f}}\ \$forget gate (8)
$\mathbf{o}_{t}\$$=\ \sigma\left(\tilde{\mathbf{o}}_{t}\right)\ ,$ $\tilde{\mathbf{o}}_{t}\$$=\ \bm{W}_{\mathbf{o}}\ \bm{x}_{t}\ +\ \bm{R}_{\mathbf{o}}\ \bm{h}_{t-1}\ +\ \bm{b}_{\mathbf{o}}\ \$$\text{output gate},$ (9)
@@
Context-free (CF) Push-down automaton (PDA) ++ infinite stack
Context-sensitive (CS) Linear bounded automaton (LBA) ++ bounded tape
Recursively enumerable (RE) Turing machine (TM) ++ infinite tape